OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line 3x + y = 7 and passes through the point (6, −1). y = one thirdx − 3 y = one thirdx + 17 y = −3x − 3 y = −3x + 17

OpenStudy (anonymous):

Ok..Two lines are perpendicular if their slopes are negative reciprocals of each other

OpenStudy (anonymous):

3x+y=7 can be written as y=-3x+7...so the slope of this line is -3 and the slope of our perpendicular line is -(1/-3)=1/3.

OpenStudy (anonymous):

it passes through (6,-1) so (y-y1)=m(x-x1) --->(y--1)=1/3(x-6) =y+1=(1/3)x-2

OpenStudy (anonymous):

y=(1/3)x-3

OpenStudy (anonymous):

@BSANDSTORM